OpenStudy (anonymous):

Given the point (2, 3) and the slope of 4, find y when x = 22. (1 point) 78 83 88 91

OpenStudy (anonymous):

y - 3 = 4 * (x - 2) plug in x = 22 to get the answer.

OpenStudy (anonymous):

88?

OpenStudy (anonymous):

nah, first add 3 to both sides then plug in the 22 in the parenthesis to solve

OpenStudy (anonymous):

you should end up with y = 7 (20)

OpenStudy (anonymous):

you should be able to take it from here lol